Integrated circuits are manufactured in vacuum chambers in which the air pressure is 8.00×10−11 mmofHgmmofHg .

You may want to review (Pages 549 - 550) .

Part A

What is the number density? Assume TTT = 22.0 ∘C∘C.

SubmitRequest Answer

Part B

What is the mean free path of a molecule?
Answer #1

Part A

first, convert the given pressure to Pa

P = 8e-11 * 101325 / 760

P = 1.0665e-8 Pa

Now,

PV = nRT

n / V = P / RT

n/V = 1.0665e-8 / 8.314 * 295

n/V = 4.348e-12 moles / m3

Now,

1 mole = 6.022e23 molecules

so,

number density = 6.022e23 * 4.348e-12

number density = 2.62e12 / m3

-----------------------------------------------------------

part B

mean free path = 1/(4 * sqrt(2) * Pi * number density * (r2))

where r = 1e-10 m

so,

mean free path = 1/(4 * sqrt(2)*Pi * 2.62e12* 1e-102))

mean free path = 2.15e6 m